Ilya and Anya each can run at a speed of 7.70 mph and walk at a speed of 3.30 mph . They set off together on a route of length 5.00 miles . Anya walks half of the distance and runs the other half, while Ilya walks half of the time and runs the other half.
Answer #1

Anya's time, t = (2.5/7.7) + (2.5/3.30) = 1.08 hrs

Average speed, v = 5/1.08 = 4.62 mph

Ilya's average speed, v = (7.7 + 3.3)/2 = 5.5 mph

Time, t = 5/5.5 = 0.91 hrs

Comment in case any doubt please rate my answer ....